OpenStudy (acespeedfighter):

1. (07.01 MC) Part A: Explain why x = 3 makes 4x − 1 ≤ 11 true but not 4x − 1 < 11. (5 points) Part B: What value from the set {4, 5, 6, 7, 8} makes the equation 8x + 1 = 57 true? Show your work. (5 points)

OpenStudy (oswaldmurphy):

You using flvs?

OpenStudy (welshfella):

replace x by 3 and see what you get 4x - 1 ≤ 11 4(3) - 1 = what ?

OpenStudy (acespeedfighter):

i dont know

OpenStudy (welshfella):

4 times 3 = ?

OpenStudy (acespeedfighter):

12

OpenStudy (welshfella):

now subtract 1 from 12 12 - 1 = 11 right/

OpenStudy (welshfella):

4(3) - 1 = 12 - 1 = 11

OpenStudy (acespeedfighter):

yes

OpenStudy (welshfella):

so that makes it true because ≤ 11 means greater than OR equal to 11

OpenStudy (welshfella):

now plug in x = 3 into the other inequality 4x - 1 < 11 4(3) - 1 = 11 so is 4x - 11 < 11 true or false when x = 3?

OpenStudy (acespeedfighter):

true

OpenStudy (welshfella):

?? what does 4x - 1 < 11 mean in words?

OpenStudy (acespeedfighter):

oh sorry i meant to say false

OpenStudy (welshfella):

< means what?

OpenStudy (acespeedfighter):

greater than

OpenStudy (welshfella):

sorry I made a mistake in an earlier post ≤ means less than or equal to not greater than or equal to

OpenStudy (welshfella):

NO < means less than

OpenStudy (acespeedfighter):

oh ok

OpenStudy (welshfella):

so 4(3) - 1 = 11 when x = 3 Now is 11 less than 11?

OpenStudy (acespeedfighter):

yea its greater than

OpenStudy (welshfella):

Of course not so for x = 3, 4x- 1 <11 is false

OpenStudy (acespeedfighter):

k

OpenStudy (welshfella):

Do you see why its false? beacuse when x = 3, 4x - 1 = 11 but the inequality is 4x - 1 < 11

OpenStudy (acespeedfighter):

yea

OpenStudy (welshfella):

For part B one way to do this is to solve the equation (find what value of x fits the equation) 8x + 1 = 57 can you do that?

OpenStudy (acespeedfighter):

nope

OpenStudy (welshfella):

OK well the other way is to plug in each of the numbers in the given set and see which one fits

OpenStudy (acespeedfighter):

8

OpenStudy (welshfella):

try x = 4; 8*4 + 1 = 32 + 1 = 33 - we need it to work out to 57 not 33 . So 4 is not the answer. Neither is 8 try the other numbers

OpenStudy (welshfella):

what is 8*5 + 1 - do the multiply first then add the 1

OpenStudy (welshfella):

you can do it Multiply 8 by 5 then add 1